Consider this short run production function, similar to many we discussed in class:

Q = (-.05)*L3 + (2)*L2



What is the value of average product (AP) when you employ 30 workers?
1
Expert's answer
2020-03-27T10:35:42-0400

Average product is: "AP = Q\/L = (-0.05L^3 + 2L^2)\/L = -0.05L^2 + 2L = -0.05*30^2 + 2*30 = 15."